a company is criticized because only 18 of 50 executive at a local branch are women. the company explaines thay althouhht this proportion is lower than it might wish it is not a suprising value given that only 43% of all its employees are women. the company has more than 500 executives worldwide. test an appropriate hypothesis and state the conclusion
Answer #1

Test statistic 2² - P-P La And from the given data = = 186 = 0.36 P=0.43 Ple 0-36-0.43 2* = 9030 - 0.43(0.53) V so z*-09998 E